# Self-Enhancing Video Data Management System for Compositional Events with Large Language Models

### v1

@@ 6.1. End-to-end performance
We first evaluate the end-to-end performance of VOCAL-UDF and baseline systems by varying the number of missing UDFs, from zero (i.e., all required UDFs are available) to three missing UDFs. In each experiment with missing UDFs, for each query separately, all base UDFs are available, but supplemental UDFs are randomly removed. Thirty distinct queries are each executed three times.
@@
VOCAL-UDF achieves high F1 scores even when UDFs are missing. Figure 5 shows the results. When all required UDFs are available, all systems perform similarly. For CLEVRER, all predefined UDFs are “perfect” since they are the same as the ones used for generating the ground-truth scene graph annotations, so F1 scores with no missing UDFs are 1.0. In contrast, CityFlow-NL and Charades have predefined UDFs that include less accurate ML models, leading to lower F1 scores compared to CLEVRER with no missing UDFs. As expected, VisProg and EQUI-VOCAL see significant drops in F1 scores when required UDFs are missing because they cannot generate new UDFs. The F1 scores decrease further as more UDFs are missing. VisProg performs worse than EQUI-VOCAL because the former is more likely to include unavailable UDFs in the generated program, rendering it non-executable. VOCAL-UDF mitigates F1 score degradation by generating new UDFs as needed. However, performance drops more for Charades, since its missing UDFs involve complex semantic relationships that are more difficult to generate with high quality.
In the following subsections, we drill down into the performance of each component of VOCAL-UDF.
6.2. Proposing UDFs
@@
Translating NL queries to DSL is not a contribution of this paper, but that step affects the end-to-end performance, so we report it. For this experiment, all UDFs are available, and we ask the LLM to translate the same queries used in the end-to-end experiment, running each query three times. We measure the F1 score of the generated queries over the dataset. About 80% of the time, the translated DSL returns the exact video segments intended by the NL queries (76% for CLEVRER, 89% for CITIFLOW-NL, and 84% for CHARADES), which increases to 93% if we consider an F1 score 0.98 (93% for CLEVRER, 96% for CITIFLOW-NL, and 97% for CHARADES). Common translation errors include missing and redundant predicates, misuse of UDFs with similar names, and incorrect temporal ordering of predicates.
@@
| CLEVRER | 540 | 456 | 23 | 107 |
| CityFlow-NL | 270 | 267 | 4 | 7 |
| Charades | 270 | 235 | 9 | 44 |
Next, we evaluate how well VOCAL-UDF proposes UDFs, considering the same set-up as in the end-to-end evaluation above. For the experiments with missing UDFs, to determine the correctness of a proposed UDF, we compare its name to the list of UDF names that are included in the target query but not in the list of available UDFs. When comparing the name, we manually inspect the results to account for synonyms (e.g., “color_yellow” vs. “yellow”, “in” vs. “inside”). As shown in Table 2, VOCAL-UDF proposes fewer UDFs than expected and has more false negatives (FNs) than false positives (FPs), because some of the new UDFs are equivalent to the available UDFs (e.g., “above(o1, o2)” is equivalent to “beneath(o2, o1)”). Specifically, there are 68, 5, and 20 missed UDFs (FNs), respectively, in the three datasets due to equivalent UDFs being available. However, this is beneficial as it avoids the unnecessary cost of generating new UDFs when equivalent ones are already available. For Charades, 13 FNs are for “holding”, which VOCAL-UDF decides to approximate using other available UDFs (e.g., “touching”). Most false positives arise from proposing compositional UDFs. For instance, VOCAL-UDF might propose “location_bottom_left” to cover both “bottom” and “left” simultaneously, or “behind_and_near” for “behind” and “near”.
6.3. Program-based UDFs generation
@@
| CLEVRER | 0.982 | 0.444 | 0.518 | 0.000 |
| CityFlow-NL | 1.000 | 0.679 | 0.359 | 0.142 |
| Charades | 0.998 | 0.080 | 0.172 | 0.009 |
We now evaluate the performance of program-based UDFs. We consider all correctly proposed program-based UDFs across the 270 experiments (3 datasets x 30 queries x 3 runs). We classify UDFs into two categories, one where at least one of their best implementations is program-based, and another where the best is not program-based.
Table 3 shows the median F1 score of the best generated program for each UDF (best UDFs) and the median F1 score of all generated programs for each UDF (all UDFs). When the best UDF generation is program-based, VOCAL-UDF demonstrates the ability to produce high-quality programs. However, not all generated UDF candidates are high quality; for instance, the median F1 score of all program-based UDFs for the Charades dataset is only . Therefore, VOCAL-UDF must carefully select the best candidate during the UDF selection phase. When the best UDF is not program-based, the F1 scores of generated programs decrease significantly, indicating that certain relationships or attributes may not be well-suited for program-based UDFs.
@@
| CLEVRER | 147 | 63 | 57 | 27 | 205 | 146 | 104 | 50 |
| CityFlow-NL | 24 | 9 | 15 | — | 241 | 215 | 77 | — |
| Charades | 72 | 10 | 62 | — | 200 | 35 | 140 | — |
As described in Section 4.2.1, VOCAL-UDF supports program generation with various enhanced features. Thus, we evaluate how often VOCAL-UDF uses these features in its program generation. We categorize program-based UDFs into four types: “all”, where every generated program fits this category; “reuse”, which utilizes results from existing UDFs; “param”, which takes numeric parameters as function inputs; and “pixel”, which accepts frame pixels as function inputs. Table 4 counts the number of the best program generations for each UDF in each category. The results show that best generated programs exhibit a variety of types, suggesting that numeric parameters, reusing results of existing UDFs, and pixels input are all important features for generating high-quality program-based UDFs.
6.4. Distilled-model UDFs generation
@@
| CLEVRER | 0.844 | 0.521 | 0.472 | 0.664 |
| CityFlow-NL | 0.653 | 0.282 | 0.695 | 0.668 |
| Charades | 0.204 | 0.024 | 0.362 | 0.721 |
Next, and using the same method as above, we evaluate the performance of distilled-model UDFs. UDFs are classified into two categories, one where at least one of their best implementations is distilled-model, and another where none are distilled-model. Table 5 shows the median F1 score for the distilled-model generation for each UDF and the median F1 score for the dummy generation for each UDF as the baseline. When the best UDF generation is distilled model, VOCAL-UDF shows a significant improvement in F1 scores over the baseline. When the best UDF is not distilled model, the F1 scores of distilled models are comparable or worse than the baseline, indicating that certain relationships or attributes may not be suitable for distilled-model UDFs. For the CityFlow-NL and Charades datasets, the F1 scores are higher compared to UDFs whose best implementations are distilled models, likely because the latter are often complex semantic relationships and attributes that are more difficult to classify.
@@
| suv | 0.782 | holding | 0.563 | sitting on | 0.781 |
| white | 0.883 | standing on | 0.868 | covered by | 0.717 |
| grey | 0.762 | carrying | 0.718 | eating | 0.547 |
| van | 0.872 | wiping | 0.671 | touching | 0.536 |
| sedan | 0.797 | leaning on | 0.756 | wearing | 0.820 |
| black | 0.784 | drinking from | 0.574 | lying on | 0.804 |
| red | 0.801 | writing on | 0.803 | above | 0.356 |
| blue | 0.874 | in front of | 0.643 | beneath | 0.310 |
| pickup truck | 0.912 | behind | 0.606 | in | 0.483 |
We further examine the VLM’s labeling quality during distilled-model UDF generation. We conduct this experiment on the CityFlow-NL and Charades datasets. We evaluate the labeling quality across 9 attributes in CityFlow-NL and 18 relationships111“have it on the back” and “twisting” are removed due to insufficient number of positives. in Charades. For each attribute or relationship, we use GPT-4V to label a randomly-constructed, balanced set of 500 samples. Table 6 shows the average F1 score of each concept over three runs. GPT-4V generally achieves high labeling quality for a wide range of concepts. Of the 27 concepts evaluated, nine achieve F1 scores of at least 0.8, and 17 attain F1 scores of at least 0.7. However, GPT-4V struggles to label the five spatial relationships, suggesting that VLMs like GPT-4V are still limited in spatial reasoning.
6.5. Choosing between program-based and distilled-model UDFs
| Dataset | best “dummy” | best “dummy” | |||
| No. | both | llm | No. | both | |
| CLEVRER | 233 | 212 (91%) | 164 (70%) | 9 | 5 (56%) |
| CityFlow-NL | 174 | 144 (83%) | 114 (66%) | 2 | 1 (50%) |
| Charades | 123 | 92 (75%) | 76 (62%) | 41 | 28 (68%) |
VOCAL-UDF supports four UDF generation strategies, as described in Section 4.4.3. We evaluate how well “both” and “llm” strategies can correctly determine the best type of a UDF, where “both” generates both types of UDFs and then selects the best one, and “llm” asks the LLM to decide ahead of time the type to generate. In cases where multiple UDF candidates with different types achieve the same highest F1 score, any of these types is considered correct. We analyze the results of the end-to-end experiment with the largest number of new UDFs, which uses the “both” strategy, and we only consider UDFs that are correctly proposed (Table 2).
Table 7 shows the number of correctly selected UDF types. The results are divided into two categories: one where at least one of the best UDF candidates is not of the “dummy” type, and another where the best type is “dummy”. When the best UDF type is not “dummy”, VOCAL-UDF correctly selects the UDF type at least 75% of the time with the “both” strategy, reaching up to a 91% correctness rate on the CLEVRER dataset. Using “llm” results in a lower correctness rate than using “both”, but still achieves at least 62%. When the best UDF type is “dummy”, VOCAL-UDF can also select the correct UDF type 50% to 68% of the time.
| Dataset | program | model | dummy |
| CLEVRER | 172 (71%) | 55 (23%) | 15 (6%) |
| CityFlow-NL | 100 (57%) | 64 (36%) | 12 (7%) |
| Charades | 79 (48%) | 34 (21%) | 51 (31%) |
Table 8 presents the distribution of UDF types selected by VOCAL-UDF using the “both” strategy. To handle queries encompassing a wide range of semantic concepts, VOCAL-UDF generates and selects UDFs of various types, underscoring the necessity for VOCAL-UDF to support two different types of UDFs. Notably, VOCAL-UDF selects more dummy UDFs (51 instances, 31%) on the Charades dataset compared to the other two datasets. Among them, 19 instances are “behind”, which involves reasoning about the distance of objects from the camera. VOCAL-UDF struggles with this task due to the lack of access to depth maps of the frames. Additionally, 18 instances are “in”, which our rule-based predictor identifies by checking if the bounding boxes of two objects overlap. This method results in a significant number of object pairs being classified as having this relationship, thereby allowing even the dummy UDFs to perform exceptionally well in these cases.
@@ 6.6. UDF selection
| Dataset | No. | best | 80% of best |
| CLEVRER | 242 | 141 (58%) | 221 (91%) |
| CityFlow-NL | 176 | 130 (74%) | 148 (84%) |
| Charades | 164 | 110 (67%) | 137 (84%) |
Finally, we evaluate the performance of UDF selection. In Table 9, the “best” column shows the number of UDFs selected by VOCAL-UDF that achieve the highest F1 score among all UDF candidates, while the “80% of best” column represents the number of selected UDFs that attain at least 80% of the best F1 score. Active learning helps VOCAL-UDF select better-performing UDFs from candidates with a labeling budget as low as 20. Although VOCAL-UDF does not always pick the best UDF due to similar scores for many candidate UDFs, it is still able to select a good UDF implementation at least 84% of the time (with an F1 score of at least 80% of the best implementation).

### v2

@@ 6.1. End-to-end performance
6.1.1. F1 score, precision, and recall.
@@
We first evaluate the end-to-end performance of VOCAL-UDF and baseline systems by varying the number of missing UDFs from zero (i.e., all required UDFs are available) to three. In each experiment, all base UDFs are available, but supplemental UDFs are randomly removed. We execute thirty distinct queries three times. Figure 5 presents the F1 scores, precision, and recall. VOCAL-UDF maintains high performance even when UDFs are missing. When all required UDFs are available, all systems perform similarly. For CLEVRER, F1 scores with no missing UDFs reach 1.0 since predefined UDFs are also used to generate ground-truth scene graph annotations. However, CityFlow-NL and Charades contain less accurate ML models in their UDFs, leading to lower F1 scores. When UDFs are missing, VisProg and EQUI-VOCAL experience significant F1 score drops, as they cannot generate new UDFs, with VisProg performing worse due to its higher likelihood of including unavailable UDFs. VOCAL-UDF mitigates F1 score degradation by generating new UDFs as needed, though performance drops more on Charades as its missing UDFs involve complex, harder-to-generate semantic relationships. Interestingly, while EQUI-VOCAL achieves higher recall than VOCAL-UDF on CLEVRER and CityFlow-NL when UDFs are missing, it does so at the expense of substantially lower precision, for an overall lower F1 score.
6.1.2. Comparing with direct LLM methods
To understand the performance gap between VOCAL-UDF and using an LLM directly, we evaluate both approaches on CLEVRER. We use the same queries from the end-to-end experiments, but under a simplified setting due to the high expense of GPT-4o. We select 500 videos and downsample each by 75%, retaining every fourth frame. We use the first 10 queries as evaluated in Section 6.1.1 but remove any duration constraints due to the downsampling. We compare VOCAL-UDF to a vanilla LLM method in which we provide GPT-4o with a sequence of video frames and ask with the prompt: “Examine the sequence of frames from a video and determine if the following event occurs? ¡user_query¿. Answer with ‘yes’ or ‘no’.” We use OpenAI’s Batch API that offers 50% lower cost. Figure 6(a) shows that VOCAL-UDF achieves significantly higher F1 scores than the vanilla LLM approach. The cost of the vanilla LLM method scales with the dataset size, averaging $5.17 USD per query under this simplified setting. If extended to the full 5,000 videos in the test set of CLEVRER, the estimated cost would be $207 USD per query.
As an alternative baseline, one can use an LLM to classify new concepts instead of generating UDFs. However, as discussed later in Section 6.1.4, the cost of doing so without downsampling is prohibitively high. While downsampling reduces LLM costs, it can also degrade query performance. Figure 6(b) shows the F1 scores of executing ground-truth DSL queries (i.e., assuming perfect LLM predictions) over the simplified CLEVRER with various downsampling factors. Downsampling greatly reduces the upper bound on achievable F1 scores, with performance dropping to 0.51 at a downsampling factor of just 8.
6.1.3. Execution time.
To analyze VOCAL-UDF’s efficiency, we perform an end-to-end evaluation of execution time. Figure 7(a) breaks down the wall-clock time for each system component, with each stacked bar showing the mean execution time for queries with different numbers of UDFs that VOCAL-UDF decides to generate. VOCAL-UDF may occasionally decide to generate extra, unnecessary UDFs. For clarity, we omit CLEVRER with four generated UDFs and CityFlow-NL with three generated UDFs from Figure 7, as each has only one data point—arising from the LLM’s inaccurate decision—and does not reflect an average execution time for that number of generated UDFs. The average execution time is under 100 seconds when all required UDFs are available and increases proportionally with the number of generated UDFs. In CLEVRER, query execution takes longer when new UDFs need to be created, requiring several minutes to process the full 14-hour video dataset as VOCAL-UDF materializes new UDF outputs. Subsequent queries using the same UDFs, however, are as fast as the case with no missing UDFs. CityFlow-NL and Charades are faster due to smaller datasets and the exclusion of frame pixels in program-based UDF generation, which eliminates the need for video loading and reduces the program execution cost. For CityFlow-NL and Charades, UDF generation dominates the execution time, as the number of VLM-annotated frames per UDF increases to 500, with active learning applied after labeling every 100 samples. UDF selection requires non-negligible time across all datasets, involving tasks such as frame loading, feature extraction, program execution, and model prediction.
Figure 7(b) provides a detailed breakdown of wall-clock time for the UDF generation process. Since VOCAL-UDF asynchronously parallelizes calls to the LLM, we measure the time breakdown as follows. All categories except for “other” capture the synchronous portions of each sub-step (i.e., excluding LLM calls), while “other” represents the remaining observed time, which includes the time spent explicitly waiting on asynchronous LLM calls. Program-based UDF generation is labeled as “program,” while distilled-model UDF generation is further divided into sub-steps (“init”, “data loading”, etc.). Distilled-model UDF generation takes significantly longer than program-based UDFs, but no single sub-step consistently dominates across all cases. Specifically, training distilled models takes the longest in CLEVRER and CityFlow-NL, whereas active learning takes the longest in Charades due to a larger dataset. The active learning step could be accelerated by performing it on a sample of the unlabeled dataset. The results also suggest that LLM calls do not dominate UDF generation time, thanks to asynchronous execution. For example, during data labeling, VOCAL-UDF submits batches of 100 asynchronous tasks to maximize parallelization.
6.1.4. Monetary cost.
To understand the query execution cost, we consider both the LLM API invocation cost and the compute resource cost. Figure 7(c) shows the mean monetary cost of OpenAI LLM calls per query. VOCAL-UDF costs less than $1.0 USD per query, with data labeling for distilled-model UDFs being the primary cost factor. Moreover, GPT-4o can be replaced by open-source models. For instance, Amazon Bedrock offers Llama 3.1 70B for a price of $0.72 USD per 1M tokens (aws, 2024a), which is cheaper for input and cheaper for output tokens than GPT-4o. Other open-source models and hosting options (lla, 2024) are available to further reduce costs, with a potential trade-off in query F1 scores. Compute resource costs depend on deployment scenarios. For example, on a g4dn.12xlarge EC2 spot instance, the estimated average resource cost of running a query with the largest number of missing UDFs is $0.7 USD for CLEVRER, $0.4 for CityFlow-NL, and $0.5 for Charades (aws, 2024b). Consequently, the total per-query cost for each dataset remains under $1.5.
| Method | GPT-4o | Llama 3.1 70B |
| VOCAL-UDF | $0.4 | $0.03–0.11 |
| Vanilla LLM | $207 | $59 |
| LLM for concepts | $1504 | $430 |
Table 3 compares the estimated LLM costs of VOCAL-UDF, “vanilla LLM”, and “LLM for concepts” (as described in Section 4.5) using both GPT-4o and Llama 3.1 70B. Due to the high expense of the baselines, we estimated costs rather than running full experiments. We estimate the cost of “LLM for concepts” as follows: Each CLEVRER video contains an average of 544 objects and 1,933 object pairs. Each LLM call requires 300 input tokens (255 for the image and 45 for the task) and one output token (“yes” or “no”). We first run the query with missing UDFs removed, then invoke the LLM only on the matching videos to minimize costs. As shown in Table 3, VOCAL-UDF is much more cost-effective than direct LLM baselines. While further optimizations—e.g., evaluating immutable attributes on a single frame and propagating them to other frames for the same object—could reduce the cost of “LLM for concepts”, the cost remains proportional to the dataset size, in contrast to VOCAL-UDF.
6.2. UDF proposal
| Dataset | F1 = 1 | F1 0.98 |
| CLEVRER | 79% | 100% |
| CityFlow-NL | 92% | 99% |
| Charades | 83% | 97% |
@@ 6.2. UDF proposal
Translating NL queries to DSL is not a contribution of this paper, but we report it as it affects the end-to-end performance. For this experiment, all UDFs are available, and the LLM translates the same queries from the end-to-end experiment, running each three times. We measure the F1 score of the generated queries. As shown in Table 4, VOCAL-UDF can effectively parse the complex textual queries into the DSL notation, returning correct video segments at least 79% of the time, which increases to if we consider an F1 score 0.98. Common translation errors include missing and redundant predicates, misuse of UDFs with similar names, and incorrect temporal ordering of predicates.
@@
| CLEVRER | 540 | 441 | 27 | 126 |
| CityFlow-NL | 270 | 261 | 4 | 14 |
| Charades | 270 | 229 | 19 | 60 |
We evaluate VOCAL-UDF’s UDF proposal using the same set-up as the end-to-end evaluation. A proposed UDF is considered correct if its name matches one of the target query’s ground truth UDFs, with a manual check for synonyms. As shown in Table 5, VOCAL-UDF proposes fewer UDFs than expected, with more false negatives (FNs) than false positives (FPs). This occurs because some new UDFs are equivalent to available UDFs (e.g., “above(o1, o2)” vs. “beneath(o2, o1)”). However, this is beneficial as it avoids unnecessary UDF generation. Most other FNs are mainly because VOCAL-UDF decides to approximate using other available UDFs (e.g., “touching” for “holding”). Most FPs arise from proposing compositional UDFs, such as “behind_and_near” for “behind” and “near”.
6.3. UDF generation
@@
| CLEVRER | 0.995 | 0.203 | 0.664 | 0.024 |
| CityFlow-NL | 1.000 | 0.831 | 0.349 | 0.203 |
| Charades | 1.000 | 0.092 | 0.162 | 0.030 |
We now evaluate the performance of program-based UDFs. We consider all correctly proposed program-based UDFs across the 270 experiments with the largest number of new UDFs (3 datasets 30 queries 3 runs). We classify UDFs into two categories, one where at least one of their best implementations is program-based, and another where the best is not program-based. Table 6 shows the median F1 score of the best generated program for each UDF (best UDFs) and all generated programs for each UDF (all UDFs). When the best UDF generation is program-based, VOCAL-UDF consistently produces high-quality programs. However, not all generated UDF candidates are of high quality; for instance, the median F1 score of all program-based UDFs for Charades is only . Thus, VOCAL-UDF must carefully select the best candidate during the UDF selection phase. When the best UDF is not program-based, F1 scores decrease significantly, indicating that certain relationships or attributes may not be well-suited to program-based UDFs.
@@
| CLEVRER | 195 | 52 | 59 | 40 | 37 | 8 | 14 | 12 |
| CityFlow-NL | 83 | 25 | 14 | — | 91 | 90 | 16 | — |
| Charades | 70 | 32 | 15 | — | 75 | 33 | 28 | — |
As described in Section 4.2.1, VOCAL-UDF supports program generation with various enhanced features. We evaluate how often VOCAL-UDF uses these features by categorizing program-based UDFs into four types: “all” (represents all generated programs), “reuse” (utilizes results from existing UDFs), “param” (takes numeric parameters as inputs), and “pixel” (accepts frame pixels as inputs). Table 7 shows the number of best generated programs for each UDF in each category. The results indicate that the best programs span various types, highlighting the importance of enhanced features in generating high-quality program-based UDFs.
@@
| CLEVRER | 0.889 | 0.519 | 0.556 | 0.662 |
| CityFlow-NL | 0.687 | 0.289 | 0.735 | 0.664 |
| Charades | 0.724 | 0.207 | 0.302 | 0.728 |
Using the same method as above, we evaluate the performance of distilled-model UDFs. UDFs are classified into two categories, one where at least one of their best implementations is distilled-model, and another where none are distilled-model. Table 8 shows the median F1 score for both distilled-model generation and dummy generation (as the baseline) for each UDF. When the best UDF generated is distilled model, VOCAL-UDF shows a significant improvement in F1 scores over the baseline. When the best UDF is not distilled model, the F1 scores of distilled models are comparable or worse than the baseline, indicating that certain relationships or attributes may not be suitable for distilled-model UDFs. For CityFlow-NL, F1 scores are higher for UDFs whose best implementations are not distilled models, likely due to the complexity of semantic attributes that are more challenging to classify.
To access VOCAL-UDF’s ability to handle ambiguous intent in program-based UDFs, we test three concepts from CLEVRER—“behind”, “far”, and “location_bottom”. For each concept, we execute the same query with one predicate but three different interpretations. For instance, “behind” is interpreted as: (1) o1’s center above o2’s center, (2) o1’s center below o2’s center, and (3) o1’s bottom edge above o2’s bottom edge. We compare VOCAL-UDF to a baseline that uses an LLM to generate a single program without numeric hyperparameters or UDF reuse. Figure 8 shows the F1 scores for each concept across varying interpretations, showing that VOCAL-UDF can more robustly adapt to different definitions of the same concept. The baseline struggles with “behind” since it uses an interpretation based on bounding box overlap that mismatches all three ground truths. While it works well for one interpretation of the other concepts, its single-generation method fails to accommodate varying interpretations, leading to significant performance drops. One limitation of VOCAL-UDF is that a concept becomes fixed once a UDF is generated. The UDF would have to be explicitly deleted or disabled if a user wanted a different interpretation for an already-existing concept. Additionally, VOCAL-UDF currently does not handle intent ambiguity for distilled-model UDFs, but this can be extended by asking LLMs to generate diverse descriptions of the target concept and distilling multiple models.
@@
| suv | 0.843 | holding | 0.776 | sitting on | 0.897 |
| white | 0.906 | standing on | 0.846 | covered by | 0.749 |
| grey | 0.797 | carrying | 0.761 | eating | 0.737 |
| van | 0.956 | wiping | 0.685 | touching | 0.632 |
| sedan | 0.903 | leaning on | 0.816 | wearing | 0.854 |
| black | 0.819 | drinking from | 0.916 | lying on | 0.857 |
| red | 0.973 | writing on | 0.894 | above | 0.683 |
| blue | 0.904 | in front of | 0.644 | beneath | 0.561 |
| pickup truck | 0.956 | behind | 0.599 | in | 0.444 |
We further examine the VLM’s labeling quality during distilled-model UDF generation on the CityFlow-NL and Charades datasets. We evaluate nine attributes in CityFlow-NL and 18 relationships in Charades. For each attribute or relationship, we use GPT-4o to label a randomly selected, balanced set of 500 samples. Table 9 shows the average F1 score of each concept over three runs. GPT-4o generally achieves high labeling quality for a wide range of concepts: seven out of the 27 concepts achieve F1 scores of at least 0.9, and 15 attain F1 scores of at least 0.8. However, GPT-4o struggles to label the five spatial relationships, suggesting that VLMs like GPT-4o are still limited in spatial reasoning.
@@ 6.4. UDF selection
| Dataset | No. | best | 80% of best | program | model | dummy |
| CLEVRER | 232 | 163 (70%) | 218 (94%) | 183 (79%) | 37 (16%) | 12 (5%) |
| CityFlow-NL | 174 | 144 (83%) | 154 (89%) | 94 (54%) | 76 (44%) | 4 (2%) |
| Charades | 145 | 93 (64%) | 123 (85%) | 81 (56%) | 23 (16%) | 41 (28%) |
Table 10 reports the UDF selection results. The “best” column shows the number of UDFs selected by VOCAL-UDF that achieve the highest F1 score among all UDF candidates, while the “80% of best” column shows how many attain at least 80% of the best F1 score. VOCAL-UDF effectively selects better-performing UDFs from candidates, even with a labeling budget as low as 20. Although VOCAL-UDF does not always pick the best UDF due to similar scores for many candidate UDFs, it can still select a good UDF implementation at least 85% of the time (with an F1 score of at least 80% of the best implementation). Table 10 also breaks down the types of UDFs that VOCAL-UDF selects. To handle diverse semantic concepts, VOCAL-UDF generates and selects UDFs of various types, underscoring the necessity for VOCAL-UDF to support two different types of UDFs. Interestingly, VOCAL-UDF selects more dummy UDFs (41 instances, 28%) on the Charades dataset than the other datasets. Notably, 20 instances are “holding”, which is difficult to distinguish under the current labeling budget. Another 17 cases are “in”, which our rule-based predictor identifies by checking bounding box overlap. This method classifies most object pairs as having this relationship, thereby allowing even the dummy UDFs to perform exceptionally well.
| Dataset | best “dummy” | best “dummy” | ||||
| No. | both | llm (gpt-4o) | llm (gpt-4-turbo) | No. | both | |
| CLEVRER | 225 | 203 (90%) | 208 (92%) | 145 (64%) | 7 | 3 (43%) |
| CityFlow-NL | 174 | 154 (89%) | 170 (98%) | 102 (59%) | 0 | 0 (—) |
| Charades | 125 | 90 (72%) | 104 (83%) | 57 (46%) | 20 | 18 (90%) |
VOCAL-UDF supports four UDF generation strategies, as described in Section 4.4.3. We now compare the “both” and “llm” strategies in selecting the correct UDF type. When multiple candidates of different types share the highest F1 score, any of these types is considered correct. Using the same end-to-end experiment setting with the largest number of missing UDFs, Table 11 shows the number of correctly selected UDF types. The results are divided into two categories: one where the best UDF type is not “dummy” and one where it is “dummy.” When the best UDF type is not “dummy,” VOCAL-UDF correctly selects the UDF type at least 72% of the time with the “both” strategy, reaching up to 90% on CLEVRER. Interestingly, using “llm” with GPT-4o yields a higher accuracy than “both”. However, this strategy is highly sensitive to the model used; switching to an earlier GPT-4 Turbo model reduces the accuracy to 46–64%. Despite this, the “llm” strategy demonstrates potential to further improve performance while reducing latency and cost. When the best UDF type is “dummy,” VOCAL-UDF can also select the correct UDF type 43% to 90% of the time using the “both” strategy.
We study the impact of active learning and dummy UDFs to the UDF selection process, using the same end-to-end experiment setting with the largest number of new UDFs. Figure 9 compares F1 scores across three system variants: the complete system with both active learning and dummy UDFs (All), the system with random sampling and dummy UDFs (No active learning), and the system with active learning but without dummy UDFs (No dummy UDFs). Both active learning and dummy UDFs help achieve higher F1 scores. For CityFlow-NL, the best UDF implementations are easily distinguishable from candidates, so active learning does not improve F1 scores but also does not degrade system performance.
VOCAL-UDF reduces user effort by using active learning to request labels at query time rather than asking the user to provide all examples upfront. Figure 10 compares the number of samples needed under two approaches—active learning during UDF selection and random sampling before query execution—to obtain at least ten positive examples for a given UDF. We consider all generated UDFs from the end-to-end experiment and classify them as “easy” (spatial relationships and attributes with relatively balanced labels) or “hard” (other concepts with greater class imbalance). Active learning allows VOCAL-UDF to collect positive samples with reduced labeling effort, particularly for “hard” concepts. Since most object pairs exhibit “in” relationships in Charades, random sampling requires labeling only 11.5 samples on average to yield 10 positives. As a result, the overall labeling effort for “easy” concepts in Charades is similar between random sampling and active learning.